Operation Summary for August 25, 2010 - August 27, 2010
Wednesday August 25
The day shift began with Operations monitoring the Tevatron (TeV) store 8010 (173.77E30) and the Recycler stash (58.5E10), with the Antiproton Source (Pbar) stacking (9.0E10), with the Linac, Booster and the Main Injector (MI) operating normally, with MiniBooNE down due to horn work, with the NuMI experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, store 8010 aborted at 1:10 PM when four TeV RF stations tripped off. RF personnel found a bad cable and replaced it. TeV personnel conducted proton-only studies. MI personnel replaced a control module for a MI RF station (MIRF6). Recycler personnel accessed the Pelletron and Switchyard personnel accessed Enclosure B.
During the evening shift, TeV studiers worked on the orbits. Operations put the TeV into HEP shot setup at 7:30 PM. At 8:07 PM, the TeV aborted due to bad beam positions at sector E11. Operators consulted with TeV personnel to fix the problem. Operations put the TeV back into shot setup at 9:50 PM and established store 8017 at 11 PM with an average initial luminosity (AIL) of 241.7E30.
Thursday August 26
The midnight shift began with Operations monitoring store 8017 (173E30) and the stash (64E10), with Pbar stacking (32E10), and with NuMI taking beam.
During the midnight shift, at 4:30 AM, the CUB LCW isolation valves for the TeV and MI closed; investigation of the cause continued until the end of shift. LRF5 tripped off several times and caused some downtime.
During the day shift, operators search and secured the MiniBooNE MI-12B enclosure. Operations began delivering beam to MiniBooNE at 10:52 AM. Operations terminated store 8017 at 11:20 AM, and put the TeV into shot setup. The TeV quenched at the B0 low beta magnets just after the store was ramped up. TeV personnel thought the beam was to close to the edge of the aperture through B0; recovery lasted the rest of the shift. Mechanical Support personnel opened the CUB isolation valves for the TeV and MI. Operators prepared Enclosure B and the collision halls for access.
During the evening shift, Operations put the TeV into shot setup at 8:34 PM and established store 8023 at 9:52 PM with an AIL of 184.8E30. Pelletron cooling problems caused some stacking downtime due to a PLC; a reboot fixed the problem.
Friday August 27
The midnight shift began with Operations monitoring store 8023 (115.13E30) and the stash (79.3E10), with Pbar stacking (12.6E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, operators spent the majority of the shift tuning Linac, Booster, and MI.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow a MI study and tuning to increase beam intensity to MiniBooNE and NuMI.
Comments and Suggestions
|last modified 08/27/2010 email Fermilab|
|Security, Privacy, Legal|